OpenStudy (anonymous):

sin inverse[sin 6pi/7]. I'm suppose to find the exact value.

OpenStudy (anonymous):

that is just 6pi/7

OpenStudy (anonymous):

that is true for all inverse functions f inverse f(x)=x

OpenStudy (anonymous):

so the same thing applies for this: cos inverse[cos pi/2] is the the exact value pi/2 ?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thanks!

OpenStudy (anonymous):

this is what inverse functions do

OpenStudy (anonymous):

what if its negative like in this case: tan inverse[tan(-pi/4)] ?

OpenStudy (watchmath):

You need to be careful here, the inverse sine function has range (-pi/2,pi/2). So the answer can't be 6pi/7.

OpenStudy (anonymous):

would you take the absolute value of "(-pi/4)" so it would be "pi/4"

OpenStudy (anonymous):

watchmath, i'm confused.

OpenStudy (anonymous):

"f inverse f(x)=x"

OpenStudy (anonymous):

oh true, did not think of that

OpenStudy (watchmath):

What I am trying to say that \(\sin^{-1}(\sin x)=x\) only for \(x\in[-\pi/2,\pi/2]\)

OpenStudy (watchmath):

So first you need to find an angle x so that sin(x)=sin(6pi/7). And we know that sin(pi/7)=sin(6pi/7) So \(\sin^{-1}(\sin 6\pi/7)=\sin^{-1}(\sin (\pi/7))=\pi/7.\)

OpenStudy (anonymous):

so the exact value is pi/7?

OpenStudy (watchmath):

Yes For the tangent, we have \[\tan^{-1}(\tan x)=x\]only for \(x\in (-\pi/2,\pi/2)\)